Calculate the 2017 total tax for Gordon Geist, a single taxpayer without dependents and no itemized deductions. He has active income of $43,000, a short-term capital gain income of $4,400 from the sale of stock, and S6,400 from book royalties. What is Gordon's average tax rate? Gordon's total gross income for the 2017 tax year is (Round to the nearest cent) Assuming Gordon's filing status is single, his standard deduction for the 2017 tax year is (Enter the amount to the nearest dollar) Click the following link for a standard deduction table Gordon is a single taxpayer without dependents. His exemption for the 2017 tax year is SD Enter the amount to the nearest dollar. Gordon's taxable income for the 2017 tax year is (Round to the nearest cent) Gordon's total income tax due for the 2017 tax year is S (Round to the nearest cent) Click the following link for the tax rates and brackets table. EEB His 2017 average tax rate is %. (Round to the nearest integer)
Standard Deduction Amounts Filing Status 2017 Single 56,350 Married Filing Jointly or Surviving Spouse Head of Household $12,700 $9,350 Married Filing Separately Personal Exemptions in 2017: $4,050 Child Tax Credit in 2017: $1,000 per child $6,350
Tax Rates and Brackets Single If Taxable Income is: The Tax is Not over $9,325 10% of taxable income Over $9,325 but not over $37,950 $932.50 plus 15% of the excess over $9.325 $5,226.25 plus 25% of the excess over $37,950 $18,713.75 plus 28% of the excess over $91,900 $46,643.75 plus 33% of the excess over $191,650 $120,910.25 plus 35% of the excess over $416,700 $121,505.25 plus 39.6% of excess over $418,400 Over $37,950 but not over $91,900 Over $91,900 but not over $191,650 Over $191,650 but not over $416,700 Over $416,700 but not over $418,400 Over $418,400 Married Filing Joint Returns and Surviving Spouses The Tax is: If Taxable Income is: Not over $18,650 10% of taxable income Over $18,650 but not over $75,900 $1,865 plus 15% of the excess over $18,650 $10,452.50 plus 25% of the excess over $75,900 $29,752.50 plus 28% of the excess over $153,100 $52,222.50 plus 33% of the excess over $233,350 $112,728 plus 35% of the excess over $416,700 $131,628 plus 39.6% of the excess over $470,700 Over $75,900 but not over $153,100 Over $153,100 but not over $233,350 Over $233,350 but not over $416,700 Over $416,700 but not over $470,700 Over $470,700
Answer #1

1). Total gross income = active income + capital gains + book royalty = 43,000 + 4,400 + 6,400 = 53,800.00

2). Standard deduction is 6,350

3). He will get one exemption of 4,050

4). Taxable income = total gross income - standard deduction - exemption =

53,800 - 6,350 - 4,050 = 43,400.00

5). Income tax due:

Over 37,950 but not over 91,900 - 5,226.25 plus 25% of the excess over 37,950

5,226.25 + 25%*(43,400 - 37,950) = 6,588.75

6). Average tax rate = 6,588.75/43,400 = 15.18% or 16%